### v1

# MaskINT: Video Editing via Interpolative Non-autoregressive Masked Transformers

Abstract
Recent advances in generative AI have significantly enhanced image and video editing, particularly in the context of text prompt control. State-of-the-art approaches predominantly rely on diffusion models to accomplish these tasks. However, the computational demands of diffusion-based methods are substantial, often necessitating large-scale paired datasets for training, and therefore challenging the deployment in practical applications. This study addresses this challenge by breaking down the text-based video editing process into two separate stages. In the first stage, we leverage an existing text-to-image diffusion model to simultaneously edit a few keyframes without additional fine-tuning. In the second stage, we introduce an efficient model called MaskINT, which is built on non-autoregressive masked generative transformers and specializes in frame interpolation between the keyframes, benefiting from structural guidance provided by intermediate frames. Our comprehensive set of experiments illustrates the efficacy and efficiency of MaskINT when compared to other diffusion-based methodologies. This research offers a practical solution for text-based video editing and showcases the potential of non-autoregressive masked generative transformers in this domain.
Figure 1: Examples of video editing with MaskINT.
1 Introduction
Text-based video editing, which aims to modify a video’s content or style in accordance with a provided text description while preserving the motion and semantic layout, plays an important role in a wide range of applications, including advertisement, live streaming, and the movie industry, etc. This challenging task requires that edited video frames not only match the given text prompt but also ensure the consistency across all video frames.
Recently, numerous studies have showcased the impressive capabilities of diffusion models [18] in the domain of text-to-image and text-to-video generation [37, 38, 3, 40]. Later on, built upon Stable Diffusion (SD) [37], several works have achieved remarkable success in the realm of text-based image editing [52, 4, 43]. When extending to text-based video editing, existing works can be mainly divided into two ways: One is to train diffusion models with temporal modules on paired text-video datasets [19, 13]. However, due to the lack of extensive text-to-video datasets, these works typically struggle to achieve the same level of editing expertise seen in the realm of image editing. The other involves leveraging a pre-trained text-to-image diffusion models in a training-free techniques [15, 34, 25, 53]. These works usually extend the self-attention across all frames to achieve an overall temporal consistency. However, this attention-based temporal constraint remains implicit and suboptimal. Moreover, while diffusion-based techniques are capable of producing high-fidelity videos, the use of diffusion models to produce all video frames proves to be a highly time-consuming process. The integration of the global attention across all frames in these video editing methods further extends the processing time, rendering them less practical for real-world applications.
Meanwhile, studies indicate that non-autoregressive masked generative transformers [7, 8, 51, 16] can attain similar levels of performance in generating images or videos compared to diffusion-based methods, while bring significant efficiency [8]. These works first tokenize image or videos into a sequence of discrete tokens [12], and then train transformers [44, 10] with masked token modeling to predict these tokens. During the inference time, they employ non-autoregressive decoding, which generates all tokens in parallel and iteratively refine predictions in a few steps. Nonetheless, extending these techniques to perform global editing tasks, such as stylization, presents a formidable challenge. This type of task requires the replacement of nearly all tokens with new ones, as opposed to merely modifying tokens within a localized region.
In this paper, we disentangle text-based video editing into two separate stages. In the first stage, we utilize existing text-based image editing models to jointly edit only two keyframes (i.e., the initial and last frames) from the video, guided by the provided text prompt. In the second stage, we propose a novel Interpolative Non-autoregressive generative Transformers (MaskINT), which performs structure-aware frame interpolation by leveraging both the color information of the initial and final frames, as well as structural cues like edge or depth maps from intermediate frames. Through disentanglement of keyframe editing and frame interpolation into separate stages, our pipeline eliminates the requirement for paired video datasets during training, thereby enabling us to train the MaskINT using video-only datasets. Furthermore, thanks to its non-autoregressive decoding, MaskINT significantly accelerates the generation of intermediate frames compared to using diffusion models for this purpose. We show that our method balances the trade off between quality and efficiency, offering comparable performance with existing diffusion methods yet taking much less time in generation.
Our major contributions are summarized as follows:
  • We propose to disentangle the text-based video editing into a two stage pipeline, that involves keyframes joint editing using existing image diffusion model and structure-aware frame interpolation with masked generative transformers trained on video only datasets.
  • We propose MaskINT to perform structure-aware frame interpolation, which is the pioneer work that explicitly introduces structure control into non-autoregressive generative transformers.
  • Experimental results demonstrate that our method achieves comparable performance with diffusion methods in terms of temporal consistency and alignment with text prompts, while providing 5-7 times faster inference times.
2 Related Work
Generative Transformers.
Following GPT [5], many pioneer works [12, 50, 26, 14, 20] tokenize images/videos into discrete tokens, and train Autoregressive Generative Transformers to perform image/video generation, where tokens are generated sequentially based on previous output. However, these autoregressive methods become exceedingly time-consuming when the length of the token sequence increases. Recently, Non-autoregressive Generative Transformers, capable of simultaneously generating all tokens in parallel, have emerged as efficient solutions [7, 51]. Specifically, MaskGiT [7] first shows the capability and efficiency of this technique in image generation. It can be seamlessly extended to tasks like inpainting and extrapolation by applying various initial mask constraints. Muse [8] achieves state-of-the-art performance in text-to-image generation by training on large-scale text-image datasets and brings significantly efficiency improvement. StyleDrop [41] further finetunes Muse with human feedback to perform text-to-image generation guided with a reference style image. Furthermore, MaskSketch [1] introduces implicit structural guidance into MaskGiT by calculating the similarity of attention maps in the sampling step. Nevertheless, this implicit structure condition is suboptimal.
In video generation, MaskViT [16] employ 2D tokenizer and trains a bidirectional window transformer to perform frame prediction. Phenaki [45] trains a masked transformer to generate short video clips condition on text prompt and extends it to arbitrary long video with different prompts in an autoregressive way. MAGVIT [51] utilizes 3D tokenizer to quantize videos and trains a single model to perform multiple video generation tasks such as inpainting, outpainting, frame interpolation, etc. However, to the best of our knowledge, there is currently no existing literature in the field of text-based video editing utilizing masked generative transformers. Besides, there is a notable absence of research that delves into explicit structural control within this area.
Diffusion Models in Image Editing.
Leveraging the advancements of SD [37], numerous studies have achieved significant success in the field of text-based image editing [52, 30, 43, 24, 17, 54, 32, 29, 9]. For example, ControlNet [52], T2I-Adapter [30], and Composer [21] finetune SD with spatial condition such as depth maps and edge maps, enabling text-to-image synthesis with the same structure as the input image. The PNP [43] incorporates DDIM inversion features [42] from the input image into the text-to-image generation process alongside SD, enabling image editing without the necessity of additional training or fine-tuning. Instructpix2pix [4] trains a conditional diffusion model for text-guided image editing using synthetic paired examples, which avoid the tedious inversion. Nevertheless, employing these methods on each video frame independently often leads to inconsistencies and flickering.
Diffusion Models in Video Editing.
Recently, diffusion models also dominate the field of video generation and video editing [40, 19]. For example, Gen-1 [13] trains a video diffusion models on paired text-video datasets to generate videos with both depth map and text prompts. Meanwhile, several works utilize pre-trained image diffusion models to achieve video editing in a training-free way [47, 49, 34, 25, 53, 46, 6]. To enable a cohesive global appearance among edited frames, a common approach in these studies involves extending the attention module of SD to encompass multiple frames and conducting cross-frame attention. In detail, Text2Video-Zero [25] performs cross-frame attention of each frame on the first frame to preserve appearance consistency. ControlVideo [53] extends ControlNet with fully cross-frame attention to joint edit all frames and further improves the performance with interleaved-frame smoother. TokenFlow [15] enhance PNP [43] with extended-attention to jointly edit a few keyframes at each denoising step and propagate them throughout the video based on nearest-neighbor field. On the contrary, we only utilize existing image diffusion models to edit two keyframes, rather than all video frames.
Video Frame Interpolation (VFI).
VFI aims to generate intermediate images between a pair of frames, which can be applied to creating slow-motion videos and enhancing refresh rate. Advanced methods typically entail estimating dense motions between frames, like optical flow, and subsequently warping the provided frames to generate intermediate ones [31, 23, 39, 28, 36, 22]. However, these methods are most effective with simple or monotonous motion. Thus, they cannot be directly applied to the second stage. In our work, we perform frame interpolation by incorporating additional structural signals.
3 Preliminaries
3.1 Masked Generative Transformers
Masked generative transformers [7] follow a two-stage pipeline. In the first stage, an image is quantized into a sequence of discrete tokens via a Vector-Quantized (VQ) auto-encoder [12]. In detail, given an image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, an encoder $\mathcal{E}$ encodes it into a series of latent vectors and discretize them through a nearest neighbour look up in a codebook of quantized embeddings with size $M$. To this end, an image can be represented with a sequence of codebook’s indices $\mathbf{Z}=[z_{i}]_{i=1}^{h\times w},z_{i}\in\{1,2,...,M\}$, where $h$ and $w$ is the resolution of latent features. A decoder $\mathcal{D}$ can reconstruct the indices back to image $\mathcal{D}(\mathbf{Z})\approx\mathbf{I}$. In the second stage, a bidirectional transformer model [44] is learned with Masked Token Modeling (MTM). Specifically, during training, a random mask ratio $r\in(0,1)$ is selected and $\lfloor\gamma(r)\cdot h\times w\rfloor$ tokens in $Z$ are replaced with a special [MASK] token, where $\gamma(r)$ is a mask scheduling function [7]. We denote the corrupted sequence with masked tokens as $\mathbf{\bar{Z}}$ and conditions such as class labels or text prompt as $\mathbf{c}$. Given the training dataset $\mathbb{D}$, a BERT [10] parameterized by $\Phi$ is learned to minimize the cross-entropy loss between the predicted and the ground truth token at each masked position:
$\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\mathbf{Z}\in\mathbb{D}}\left[\sum_{\bar{z_{i}}=\texttt{[MASK]}}-\log p_{\Phi}(z_{i}|\mathbf{\bar{Z}},\mathbf{c})\right].$ (1)
During inference time, non-autoregressive decoding is applied to generate images. Specifically, given the conditions $\mathbf{c}$, all tokens are initialized as [MASK] tokens. At step $k$, all tokens are predicted in parallel while only tokens with the highest prediction scores are kept. The remaining tokens with least prediction scores are masked out and regenerated in the next iteration. The mask ratio is determined by $\gamma(\frac{k}{K})$ at step $k$, where $K$ is the total number of iteration steps.
3.2 ControlNet
Latent Diffusion Models
Denoising Diffusion Probabilistic Models (DDPM) [18] generate images through a progressive noise removal process applied to an initial Gaussian noisy image, carried out over a span of $T(>>K)$ time steps. To enable efficient high-resolution image generation, Latent Diffusion models [37] operates the diffusion process in the latent space of an autoencoder. First, an encoder $\mathcal{E^{\prime}}$ compresses an image $\mathbf{I}$ to a low-resolution latent code $x=\mathcal{E^{\prime}}(\mathbf{I})\in\mathbb{R}^{h\times w\times c}$. Second, a U-Net $\epsilon_{\theta}$ with attention modules [44] is trained to remove the noise with loss function:
$\mathcal{L}_{LDM}=\mathbb{E}_{x_{0},\epsilon\sim N(0,I),t\sim}\|\epsilon-\epsilon_{\theta}(x_{t},t,\tau)\|_{2}^{2},$ (2)
where $\tau$ is the text prompt and $x_{t}$ is the noisy latent sample at timestep $t$. Stable Diffusion is trained on datasets with billion scale text-image pairs, which serves as the foundation model of many generation tasks.
ControlNet
In practice, it’s challenging to use text prompt to describe the layout of generated image. Furthermore, ControlNet [52] is proposed to provide spatial layout conditions such as edge map, depth map, and human poses. In detail, ControlNet train the same U-Net architecture as SD and finetune it with specific conditions. We denote ControlNet as $\epsilon_{\theta}^{\prime}(x_{t},t,\tau,s)$, where $s$ is the spatial layout condition.
Figure 2: Overview of MaskINT. MaskINT disentangle the video editing task into two separate stages, i.e., keyframes joint editing and structure-aware frame interpolation.
4 Methodology
4.1 Overview
Fig. 2 shows an overview of our framework. We disentangle the video editing task into keyframe joint editing stage and structure-aware frame interpolation stage. Specifically, given a video clip with $N$ frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$, in the first stage, with the input text prompt $\mathbf{\tau}$, we simultaneously edit two keyframes, i.e., the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$, using existing image-editing model $g(\cdot)$ that requires no additional tuning. This frame based joint editing module provides high-quality coherent edited frames $\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N}=g(\mathbf{I}^{1},\mathbf{I}^{N},\mathbf{\tau})$ and is highly efficient on a pair of frame. In the second stage, we propose MaskINT to perform structure-aware frame interpolation via non-autoregressive transformers. The entire edited video frames are generated by $\{\mathbf{{}^{*}I}^{i}\}_{i=1}^{N}=f_{\Phi}(\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N},\{\mathbf{S}^{i}\}_{i=1}^{N})$, where $\mathbf{S}^{i}\in[0,1]^{H\times W\times 1}$ is the structural condition (i.e., the HED edge map [48]). Our MaskINT is trained with masked token modeling (MTM) on video only datasets, conditioning on the structural signal $\{\mathbf{S}^{i}\}_{i=1}^{N}$ as well as the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$.
4.2 Keyframes Joint Editing
To maintain the structure layout of selected keyframes, we take ControlNet [52] to edit $\mathbf{I}^{1}$ and $\mathbf{I}^{N}$ based on text prompt $\tau$ as well as their edge maps $\mathbf{S}^{1}$ and $\mathbf{S}^{N}$. However, even with the identical noise, applying ControlNet to each keyframes individually, i.e., $\epsilon_{\theta}^{\prime}(x_{t}^{1},t,\tau,\mathbf{S}^{1})$ and $\epsilon_{\theta}^{\prime}(x_{t}^{N},t,\tau,\mathbf{S}^{N})$, cannot guarantee the appearance consistency. To address this issue, following previous work [47, 15, 53], we extend the self-attention blocks to simultaneously process two keyframes. In detail, the self-attention block projects the noisy feature map $x_{t}^{j}$ of $j^{th}$ frame at time step $t$ into query $\mathbf{Q}^{j}$, key $\mathbf{K}^{j}$, value $\mathbf{V}^{j}$ in the original U-Net of SD. We extend the self-attention block to perform attention across all selected keyframes by concatenating their keys and values and calculate the attention by
$\text{Softmax}(\dfrac{\mathbf{Q}^{j}[\mathbf{K}^{1},\mathbf{K}^{N}]^{\mathbf{T}}}{\sqrt{c}})[\mathbf{V}^{1},\mathbf{V}^{N}]$ (3)
Note that although only two frames were used in Eq.3, this joint editing can seamlessly generalize to any number of frames, but at the cost of longer processing time and huge demand in resources.
4.3 MaskINT
Structural-aware embeddings.
Previous works [16, 51] demonstrate that non-autoregressive masked generative transformers can effectively perform frame prediction and interpolation tasks. However, these architectures lack explicit control of structure, making it difficult to follow the motion of original videos. In our work, we explicitly introduce structural condition of each frame into the generation process. Specifically, we tokenize both RGB frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$ and structure maps $\{\mathbf{S}^{i}\}_{i=1}^{N}$ with an off-the-shelf 2D VQ tokenizer [12]. We utilize 2D VQ rather than 3D VQ [51] to accommodate varying numbers of frames and frame rate without constraints. We denote the tokens from RGB frames as $\mathbf{{}^{c}Z}=\{^{c}z_{i}\}_{i=1}^{N\times h\times w}$ (color token) and tokens from edge maps as $\mathbf{{}^{s}Z}=\{^{s}z_{i}\}_{i=1}^{N\times h\times w}$ (structure token), where ${}^{s}z_{i}$, ${}^{c}z_{i}\in\{1,2,...,M\}$, where $M$ is the codebook size. Subsequently, two distinct embedding layers $e^{C}(\cdot)$ and $e^{S}(\cdot)$ are learned to map token indices $\mathbf{{}^{c}Z}$ and $\mathbf{{}^{s}Z}$ into their respective embedding spaces. Learnable 2D spatial positional encoding $\mathbf{P^{S}}$ and temporal positional encoding $\mathbf{P^{T}}$ are also added [2]. Thus, the input to the following transformer layers can be formulated by $\mathbf{X}=e^{c}(\mathbf{{}^{c}Z})+e^{s}(\mathbf{{}^{s}Z})+\mathbf{P^{S}}+\mathbf{P^{T}}\in\mathbb{R}^{N\times h\times w\times c}$.
Transformer with Window-Restricted Attention.
Previous masked generative transformers [7, 51] employ a pure transformer with global attention [44]. However, Given that there is no substantial motion between consecutive frames, we adopt self-attention within a restricted window to further mitigate computational overhead, following MaskViT [16]. In detail, our approach involves two distinct stages of attention. Initially, we employ spatial window attention, confining attention to tokens within a frame of dimensions $1\times h\times w$. Subsequently, we extend this to spatial-temporal window attention, which confines attention to tokens within a tube of dimensions $N\times h_{w}\times w_{w}$, where $h_{w}$ and $w_{w}$ is the window size, where $h_{w}<h$ and $w_{w}<w$, which greatly reduce the complexity. Besides, to further reduce computation of transformer, we also add a shallow convolution layers to downsample $\mathbf{X}$ before the transformer encoder layers and an upsample layer at the end.
Training.
By fully disentangling keyframes editing and frame interpolation into distinct stages, our model no longer necessitates paired videos for training. Consequently, we can train MaskINT using video only datasets. Denote the color token of $i^{th}$ RGB frame as $\mathbf{{}^{c}Z}^{i}=\{^{c}z^{i}\}_{i=1}^{h\times w}$. During the training time, we keep color tokens of the initial frame $\mathbf{{}^{c}Z}^{1}$ and last frame $\mathbf{{}^{c}Z}^{N}$, and randomly replace $[\gamma(r)\cdot(N-2)\cdot N]$ color tokens of intermediate frames with the [MASK] tokens. We denote this corrupted video color tokens as $\mathbf{\bar{{}^{c}Z}}=\{\mathbf{{}^{c}Z}^{1},\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1},\mathbf{{}^{c}Z}^{N}\}$. The structure-aware window-restricted transformer with parameters $\Theta$ is trained by
$\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\mathbf{{}^{c}Z},\mathbf{{}^{s}Z}\in\mathbb{D}}[\sum_{\bar{{}^{c}z_{i}}=\texttt{[MASK]}}-\log p_{\Theta}(^{c}z_{i}|\mathbf{\bar{{}^{c}Z}},\mathbf{{}^{s}Z})]$ (4)
Inference.
During the inference time, our MaskINT can seamlessly generalize to perform frame interpolation between the jointly edited frames, although it is only trained with regular videos. Specifically, we tokenize the the initial and last edited frames $\mathbf{{}^{*}I}^{1}$ and $\mathbf{{}^{*}I}^{N}$ from Stage 1 into color tokens $\mathbf{{}^{c}_{*}Z}^{1}$ and $\mathbf{{}^{c}_{*}Z}^{N}$, and initialize color tokens of all intermediate frames $\{\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1}\}$ with [MASK] tokens. We follow the iterative decoding in MaskGiT [7] with a total number of $K$ steps. At step $k$, we predict all color tokens in parallel and keep tokens with the highest confidence score.
5 Experiments
5.1 Settings
Implementation Details.
We train our model with $100k$ videos from ShutterStock website [^1] https://www.shutterstock.com/video. During training time, we random select a $T=16$ video clip with frame interval $1,2,4$ from each video and resize it to $384\times 672$. We utilize Retina-VQ [11] with $8$ downsample ratio, i.e., each frame has $48\times 84$ tokens. We employ Transformer-Base as our MaskINT and optimized it from scratch with the AdamW optimizer [27] for a duration of $100$ epochs. The initial learning rate is set to $41e-$ 4 and decayed with cosine schedule. During the inference time, we set the number of decoding step $K$ to $32$ and the temperature $t$ to $4.5$.
Evaluation.
Following [47], we use the selected $40$ object-centric videos of the DAVIS dataset [33], covering humans, animals, vehicles, etc. Besides, we also select $30$ unseen videos from the ShutterStock dataset. For each video, we manually design $5$ edited prompts, including object editing, background changes and style transfers. Following previous works [53, 34], we assess the quality of the generated videos using CLIP [35]. In detail, we evaluate 1) Temporal consistency, which calculates the average cosine similarity of all pairs of consecutive frames. 2) Prompt consistency, which calculates the average cosine similarity between given text prompt and all video frames. To evaluate the efficiency, we report the duration required for generating a 16-frame video clip on a single NVIDIA A6000 GPU.
Figure 3: Examples of video editing with MaskINT. Frames with red bounding box are jointly edited keyeframes.
5.2 Results
We select methods that built upon text-to-image diffusion models for comparison, including TokenFlow [15], Text-to-video zero [25], and ControlVideo [53]. We also consider apply ControlNet to each frame individually with the same initial noise as baseline. Besides, we also compare with frame interpolation with FILM [36] with the same edited keyframes.
Method DAVIS ShutterStock Time
T.C. P.C. T.C. P.C
ControlNet per frame [52] 0.9137 0.3136 0.9419 0.3040 50s
Text2Video-zero [25] 0.9642 0.3124 0.9811 0.3036 60s
TokenFlow [15] 0.9774 0.3169 0.9869 0.3133 150s
ControlVideo-edge [53] 0.9746 0.3143 0.9864 0.3032 120s
MaskINT (ours) 0.9519 0.3112 0.9714 0.3038 22s
Table 1: Quantitative comparisons on the quality of the generated videos. “T.C.” stands for “temporal consistency”, and “P.C.” stands for “prompt consistency”.
Quantitative Comparisons.
Table 1 summarize the performance of these methods on both DAVIS and ShutterStock datasets. Notably, our method achieves comparable performance with diffusion methods, in terms of both temporal consistency and prompt consistency, while brings a significant acceleration in processing speed. In detail, MaskINT is almost 5.5 times faster than ControlVideo [53], whose fully cross-frame attention is computationally extensive. Moreover, MaskINT is nearly 7 times faster than TokenFlow [15], whose DDIM inversion is time-consuming. On the contrary, our acceleration is derived from a combination of a lightweight network design and a reduced number of decoding steps in masked generative transformers.
Figure 4: Qualitative comparisons with diffusion-based methods. More examples are shown in Supplementary.
Qualitative Comparisons.
Fig. 1 and Fig. 3 show several samples of edited videos with MaskINT. Our method is capable of generating temporally consistent videos that adhere to the provided text prompts. This extends to a wide range of applications with text prompts, encompassing tasks such as stylization, background editing, foreground editing, and more. It also works well on challenging videos with substantial motion, such as jumping and running. Moreover, Fig. 4 provides a qualitative comparison of MaskINT to other baselines. Remarkably, diffusion methods [15, 25, 53] can ensure the consistency of overall appearance, but sometimes cannot maintain the consistency of detailed regions. For example, both TokenFlow [15] and Text2Video-Zero[25] exhibit noticeable artifacts in the leg region of the human subjects. ControlVideo [53] produces inconsistent hats. The potential explanation lies in the fact that these methods offer control over temporal consistency implicitly. Furthermore, FILM [36] produces videos that deviate from the original motions with the same edited keyframes. Our MaskINT consistently interpolates the intermediate frames based on the structure condition and even maintain better consistency in local regions.
Extension on Long Video Editing.
Since the non-autoregressive pipeline generates all video frames simultaneously, it’s challenging for it to edit an entire long video due to GPU memory limitation. Nevertheless, our framework can still be extended to generate long videos by dividing the long video into short clips and progressively performing frame interpolation within each clip. For instance, given a video with $60$ frames, we select the $1^{\text{st}}$, $16^{\text{th}}$, $31^{\text{st}}$, $46^{\text{th}}$, and $60^{\text{th}}$ frames as keyframes. We jointly edit these selected $5$ frames together and then perform structure-aware frame interpolation within each pair of consecutive keyframes. As shown in Fig. 5, with this design, our method can still generate consistent long videos. Besides, in this proposed extension, the generation of later frames is decoupled from the generated early frames, which differs from the autoregressive long-video generation pipeline in Phenaki [45]. Consequently, even if some early frames encounter difficulties, the generation of later frames can still proceed successfully.
Figure 5: Examples of long video generation. The number indicates the index of frame. More examples are in Supplementary.
5.3 Ablation Studies
Video Frame Interpolations
We conduct a quantitative evaluation on the performance of our structure-aware interpolation module. In this ablation study, we perform frame interpolation using the original keyframes to reconstruct the original intermediate frames, and compare the interpolated frames with the original video frames at the pixel level. We use same testing samples from DAVIS and Shutterstock datasets, employing peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and structured similarity (SSIM) as the evaluation metrics. We benchmark our method against two state-of-the-art VFI methods, namely FILM [36] and RIFE [22]. We also showcase the performance of applying VQ-GAN [12] to all video frames, serving as an upper bound for our method.
As in Table 2, our method significantly outperforms VFI methods on all evaluation metrics, with the benefit of the structure guidance from the intermediate frames. Furthermore, Fig. 6 shows qualitative comparison between video frame interpolation methods FILM [36] and MaskINT. Even when confronted with significant motion between two frames, our approach successfully reconstructs the original video, maintaining consistent motion through the aid of structural guidance. In contrast, FILM introduces undesirable artifacts, including disorted background, duplicated cat hands, and the absence of a camel’s head, etc. The major reason is that current VFI models mainly focus on generating slow-motion effects and enhancing frame rate, making them less effective in handling frames with large motions, which usually requires a better semantic understanding. Additionally, the absence of structural guidance poses a challenge for these VFI methods in accurately aligning generated videos with the original motion.
Figure 6: Qualitative comparisons on video reconstruction with original RGB frames. Frames with red bounding box are given.
Method DAVIS ShutterStock
[…]
RIFE [22] 17.31 0.5195 0.2512 20.44 0.7210 0.1533
FILM [36] 17.00 0.5011 0.2363 20.90 0.7453 0.1246
MaskINT 22.15 0.6332 0.1483 24.19 0.7616 0.1097
VQGAN (ground truth) 25.66 0.7429 0.0784 27.81 0.8327 0.0561
Table 2: Quantitative comparisons on video frame interpolation with original keyframes.
Number of of keyframes
Although our model is trained with frame interpolation by default, MaskINT can seamlessly generalize to an arbitrary number of keyframes without finetuning. We assess the impact of varying the quantity of keyframes on the generation performance. As shown in the left part of Table 3, with an increase in the number of keyframes, the model exhibits an improvement in performance. Generally, with more information, performing frame interpolation is easier. However, simultaneously editing more keyframes requires longer time due to the global attention among them.
Decoding steps
We also explore the number of decoding steps of the masked generative transformers in the second stage. The right part of Table 3 shows that more decoding steps can bring slight improvement on the temporal consistency, but requires more time. Considering the trade-off between performance and efficiency, we chose $K=32$ steps by default in all experiments.
# keyframes T.C. P.C. Time 1 0.9690 0.2984 19s 2 0.9714 0.3038 22s 3 0.9721 0.3051 26s 4 0.9728 0.3069 29s 6 0.9737 0.3035 35s
#decoding step $K$ T.C. P.C. Time 16 0.9691 0.3038 15s 32 0.9714 0.3038 22s 64 0.9719 0.3040 33s 128 0.9720 0.3041 62s
Table 3: Ablation study on ShutterStock dataset of the number of keyframes and the number of decoding steps $K$. “T.C.” stands for “temporal consistency” and “P.C.” stands for “prompt consistency”.
6 Limitation and Future Work
One potential limitation of our approach is that it still necessitates the use of diffusion-based methods for keyframe editing in the initial stage, which can still be somewhat time-consuming. As for the future work, given that our approach disentangles video editing into two distinct stages, we intend to explore the integration of token-based methods, like Muse [8], for image editing. This endeavor aims to further enhance the efficiency of the initial stage.
7 Conclusion
We propose MaskINT towards consistent and efficient video editing with text prompt. MaskINT disentangle this task into keyframes joint editing with diffusion methods and structure-aware frame interpolation with non-autoregressive masked transformers. Experimental results demonstrate that MaskINT achieves comparable performance with pure diffusion-based methods while significantly reduces the inference time. Moreover, our work demonstrates the substantial promise of non-autoregressive generative transformers within the realm of video editing.
Supplementary Material
8 Diverse Structural Conditions
Although we by default utilize HED edge map as our structure condition for both stages, our method can also employ other structural controls in both stages due to the disentanglement. In this study, we explore utilizing ControlNet with depth map to perform key frame editing, and using the depth map as the guidance to perform structure-aware frame interpolation. Additionally, we explore various combinations of these approaches. As summarized in Table 4, all of these combinations achieve the same level of performance. The performance of prompt consistency (P.C.) is determined by the specific key frame editing methods employed. For the second stage, depth control typically offers greater flexibility than HED edge control for frame interpolation. This could be the potential reason for the slightly worse performance in temporal consistency with edge-based key frame editing.
Stage1 Stage2 T.C. P.C
HED edge HED edge 0.9714 0.3038
depth map HED edge 0.9713 0.3159
HED edge depth map 0.9683 0.3035
depth map depth map 0.9719 0.3171
Table 4: Quantitative comparisons of the combination of varied structural conditions in each stage on ShutterStock. “T.C.” stands for “temporal consistency”, and “P.C.” stands for “prompt consistency”.
9 Additional Examples for Comparisons
We add more comparisons with diffusion methods in Fig. 8. Our methods can maintain the temporal consistency among variant examples like other pure-diffusion methods.
10 Additional Editing Examples
We present more video editing examples in Figure 9 to show the generalization of our method.
11 Example of Failure Cases
Since we disentangle the video editing tasks into two separate stage, the final performance of the generated video depends on the key frame editing in the first stage. In certain challenging scenarios, the attention-based key frame editing stage struggles to produce consistent frames, primarily due to the complexity of the scene or the presence of exceptionally large motion. In this case, our MaskINT can still interpolate the intermediate frames, albeit with the potential for introducing artifacts. Figure 7 show some failure cases when the first stage fails.
Figure 7: Examples of failure cases.
Figure 8: Additional Qualitative comparisons with diffusion-based methods. Frames with red bounding box are jointly edited keyeframes.
Figure 9: Additional Editing examples with MaskINT. Frames with red bounding box are jointly edited keyeframes.

### v2

@@
 State-of-the-art approaches predominantly rely on diffusion models to accomplish these tasks.
-However, the computational demands of diffusion-based methods are substantial, often necessitating large-scale paired datasets for training, and therefore challenging the deployment in practical applications.
-This study addresses this challenge by breaking down the text-based video editing process into two separate stages.
-In the first stage, we leverage an existing text-to-image diffusion model to simultaneously edit a few keyframes without additional fine-tuning.
-In the second stage, we introduce an efficient model called MaskINT, which is built on non-autoregressive masked generative transformers and specializes in frame interpolation between the keyframes, benefiting from structural guidance provided by intermediate frames.
-Our comprehensive set of experiments illustrates the efficacy and efficiency of MaskINT when compared to other diffusion-based methodologies.
+However, the computational demands of diffusion-based methods are substantial, often necessitating large-scale paired datasets for training, and therefore challenging the deployment in real applications.
+To address these issues, this paper breaks down the text-based video editing task into two stages.
+First, we leverage an pre-trained text-to-image diffusion model to simultaneously edit few keyframes in an zero-shot way.
+Second, we introduce an efficient model called MaskINT, which is built on non-autoregressive masked generative transformers and specializes in frame interpolation between the edited keyframes, using the structural guidance from intermediate frames.
+Experimental results suggest that our MaskINT achieves comparable performance with diffusion-based methodologies, while significantly improve the inference time.
 This research offers a practical solution for text-based video editing and showcases the potential of non-autoregressive masked generative transformers in this domain.
@@
 This challenging task requires that edited video frames not only match the given text prompt but also ensure the consistency across all video frames.
-Recently, numerous studies have showcased the impressive capabilities of diffusion models [18] in the domain of text-to-image and text-to-video generation [37, 38, 3, 40].
-Later on, built upon Stable Diffusion (SD) [37], several works have achieved remarkable success in the realm of text-based image editing [52, 4, 43].
-When extending to text-based video editing, existing works can be mainly divided into two ways: One is to train diffusion models with temporal modules on paired text-video datasets [19, 13].
-However, due to the lack of extensive text-to-video datasets, these works typically struggle to achieve the same level of editing expertise seen in the realm of image editing.
-The other involves leveraging a pre-trained text-to-image diffusion models in a training-free techniques [15, 34, 25, 53].
-These works usually extend the self-attention across all frames to achieve an overall temporal consistency.
-However, this attention-based temporal constraint remains implicit and suboptimal.
-Moreover, while diffusion-based techniques are capable of producing high-fidelity videos, the use of diffusion models to produce all video frames proves to be a highly time-consuming process.
-The integration of the global attention across all frames in these video editing methods further extends the processing time, rendering them less practical for real-world applications.
-Meanwhile, studies indicate that non-autoregressive masked generative transformers [7, 8, 51, 16] can attain similar levels of performance in generating images or videos compared to diffusion-based methods, while bring significant efficiency [8].
-These works first tokenize image or videos into a sequence of discrete tokens [12], and then train transformers [44, 10] with masked token modeling to predict these tokens.
-During the inference time, they employ non-autoregressive decoding, which generates all tokens in parallel and iteratively refine predictions in a few steps.
-Nonetheless, extending these techniques to perform global editing tasks, such as stylization, presents a formidable challenge.
-This type of task requires the replacement of nearly all tokens with new ones, as opposed to merely modifying tokens within a localized region.
-In this paper, we disentangle text-based video editing into two separate stages.
-In the first stage, we utilize existing text-based image editing models to jointly edit only two keyframes (i.e., the initial and last frames) from the video, guided by the provided text prompt.
-In the second stage, we propose a novel Interpolative Non-autoregressive generative Transformers (MaskINT), which performs structure-aware frame interpolation by leveraging both the color information of the initial and final frames, as well as structural cues like edge or depth maps from intermediate frames.
-Through disentanglement of keyframe editing and frame interpolation into separate stages, our pipeline eliminates the requirement for paired video datasets during training, thereby enabling us to train the MaskINT using video-only datasets.
-Furthermore, thanks to its non-autoregressive decoding, MaskINT significantly accelerates the generation of intermediate frames compared to using diffusion models for this purpose.
-We show that our method balances the trade off between quality and efficiency, offering comparable performance with existing diffusion methods yet taking much less time in generation.
+Recently, numerous studies have showcased the impressive capabilities of diffusion models [22] in image generation and editing with text prompts [44, 13, 39, 45, 61, 6, 51].
+When extending these foundation text-to-image (T2I) models to video generation and editing, a challenge is to maintain the temporal consistency.
+Existing solutions can be mainly divided into two ways: One is to finetune T2I models with additional temporal modules on paired text-video datasets [23, 17, 9, 32, 5].
+However, the lack of extensive paired text-video datasets hinders these works from attaining the same level of performance as seen in image editing.
+The other involves leveraging a T2I models in a zero-shot manner [19, 41, 29, 8, 62].
+These works find that extending the spatial self-attention across frames can produce consistent content across frames [56].
+However, extended attention can lead to inconsistent editing of details due to its implicit temporal constraints.
+Moreover, while these techniques can generate high-fidelity videos, using cumbersome diffusion models to generate all frames is time-consuming.
+The integration of the extended spatial-temporal attention further exacerbates the running time, making them less practical for real-world applications.
+This raises the question: Is it necessary to utilize tedious diffusion models to generate all video frames?
+In this paper, we propose MaskINT, an efficient prompt-based video editing framework designed to generate high-quality videos swiftly (Fig.
+1).
+MaskINT disentangles the task into two separate stages.
+In the first stage, we utilize pre-trained T2I models with extended attention to jointly edit only two keyframes (i.e., the initial and last frames) from the video clip, guided by the provided text prompt.
+In the second stage, we introduce a novel structure-aware frame interpolation module based on non-autoregressive generative Transformers [10, 11, 60], which generates all intermediate frames in parallel with structural cues and iteratively refine predictions in a few steps.
+Our framework offers several advantages: First, through the disentanglement of frame editing and interpolation, MaskINT eliminates the requirement for paired text-video datasets during training, thereby enabling us to train the MaskINT using large-scale video-only datasets.
+Second, the usage of non-autoregressive generation significantly accelerates the processing time.
+Experimental results indicate that MaskINT achieves comparable performance with pure diffusion methods in terms of temporal consistency and alignment with text prompts, while provides 5-7 times faster inference time.
+Our work not only provides a practical solution for text-based video editing in terms of the trade-off between quality and efficiency, but also highlights the potential of masked generative transformers in this domain.
 Our major contributions are summarized as follows:
-  • We propose to disentangle the text-based video editing into a two stage pipeline, that involves keyframes joint editing using existing image diffusion model and structure-aware frame interpolation with masked generative transformers trained on video only datasets.
-  • We propose MaskINT to perform structure-aware frame interpolation, which is the pioneer work that explicitly introduces structure control into non-autoregressive generative transformers.
-  • Experimental results demonstrate that our method achieves comparable performance with diffusion methods in terms of temporal consistency and alignment with text prompts, while providing 5-7 times faster inference times.
+  • We introduce MaskINT, a two-stage pipeline for text-based video editing, which involves keyframes joint editing and structure-aware frame interpolation, eliminating the need for paired text-video datasets.
+  • The proposed structure-aware frame interpolation is the pioneer work that explicitly introduces structure control into non-autoregressive generative transformers.
+  • Experimental results demonstrate that MaskINT achieves comparable performance with diffusion methods, while significantly reducing the inference time.
 2 Related Work
 Generative Transformers.
-Following GPT [5], many pioneer works [12, 50, 26, 14, 20] tokenize images/videos into discrete tokens, and train Autoregressive Generative Transformers to perform image/video generation, where tokens are generated sequentially based on previous output.
-However, these autoregressive methods become exceedingly time-consuming when the length of the token sequence increases.
-Recently, Non-autoregressive Generative Transformers, capable of simultaneously generating all tokens in parallel, have emerged as efficient solutions [7, 51].
-Specifically, MaskGiT [7] first shows the capability and efficiency of this technique in image generation.
-It can be seamlessly extended to tasks like inpainting and extrapolation by applying various initial mask constraints.
+Following GPT [7], previous works [16, 59, 31, 18, 24] first tokenize images/videos into discrete tokens, and train Autoregressive Generative Transformers to perform image/video generation, where tokens are generated sequentially based on previous outputs.
+However, these methods become exceedingly time-consuming when the length of the token sequence increases.
+Recently, Non-autoregressive Generative Transformers have emerged as efficient solutions [10, 11, 60, 20, 54].
+These works train bidirectional transformers [14] with masked token modeling (MTM).
+During inference, all discrete tokens are generated in parallel and refined iteratively in a few steps, known as non-autoregressive decoding.
+Specifically, in image generation, MaskGiT [10] first shows the capability and efficiency of non-autoregressive transformers.
+It can be seamlessly extended to tasks like inpainting and outpainting by applying various initial mask constraints.
 Muse [11] achieves state-of-the-art performance in text-to-image generation by training on large-scale text-image datasets and brings significantly efficiency improvement.
 StyleDrop [48] further finetunes Muse with human feedback to perform text-to-image generation guided with a reference style image.
-Furthermore, MaskSketch [1] introduces implicit structural guidance into MaskGiT by calculating the similarity of attention maps in the sampling step.
-Nevertheless, this implicit structure condition is suboptimal.
+MaskSketch [3] introduces implicit structural guidance into MaskGiT by calculating the similarity of attention maps in the sampling step.
 In video generation, MaskViT [20] employ 2D tokenizer and trains a bidirectional window transformer to perform frame prediction.
@@
 Diffusion Models in Image Editing.
-Leveraging the advancements of SD [37], numerous studies have achieved significant success in the field of text-based image editing [52, 30, 43, 24, 17, 54, 32, 29, 9].
-For example, ControlNet [52], T2I-Adapter [30], and Composer [21] finetune SD with spatial condition such as depth maps and edge maps, enabling text-to-image synthesis with the same structure as the input image.
-The PNP [43] incorporates DDIM inversion features [42] from the input image into the text-to-image generation process alongside SD, enabling image editing without the necessity of additional training or fine-tuning.
+Built upon the latent diffusion models (LDM) [44], numerous studies have achieved significant success in text-based image editing [61, 36, 51, 28, 21, 63, 38, 35, 12].
+For example, ControlNet [61], T2I-Adapter [36], and Composer [25] finetune LDM with spatial condition such as depth maps and edge maps, enabling text-to-image synthesis with the same structure as the input image.
+The PNP [51] incorporates DDIM inversion features [49] from the input image into the T2I generation process, enabling zero-shot image editing with pre-trained T2I models.
 InstructPix2Pix [6] trains a conditional diffusion model for text-guided image editing using synthetic paired examples, which avoid the tedious inversion.
-Nevertheless, employing these methods on each video frame independently often leads to inconsistencies and flickering.
+Nevertheless, applying these T2I models on each frame independently usually leads to inconsistencies and flickering.
 Diffusion Models in Video Editing.
-Recently, diffusion models also dominate the field of video generation and video editing [40, 19].
-For example, Gen-1 [13] trains a video diffusion models on paired text-video datasets to generate videos with both depth map and text prompts.
-Meanwhile, several works utilize pre-trained image diffusion models to achieve video editing in a training-free way [47, 49, 34, 25, 53, 46, 6].
-To enable a cohesive global appearance among edited frames, a common approach in these studies involves extending the attention module of SD to encompass multiple frames and conducting cross-frame attention.
+Diffusion models also dominate the field of video generation and video editing [47, 23, 9, 56].
+Due to the lack of billion-scale paired video datasets, these works usually finetune existing T2I foundation models [44] with additional temporal modules on paired text-video datasets [17] or overfit a single video [56].
+Meanwhile, several works explore using T2I models in a zero-shot manner for video editing [56, 58, 41, 29, 62, 55, 8].
+To enable a cohesive appearance, a common approach in these studies involves extending the spatial self-attention module to spatial-temporal attention module and conducting cross-frame attention.
 In detail, Text2Video-Zero [29] performs cross-frame attention of each frame on the first frame to preserve appearance consistency.
-ControlVideo [53] extends ControlNet with fully cross-frame attention to joint edit all frames and further improves the performance with interleaved-frame smoother.
-TokenFlow [15] enhance PNP [43] with extended-attention to jointly edit a few keyframes at each denoising step and propagate them throughout the video based on nearest-neighbor field.
-On the contrary, we only utilize existing image diffusion models to edit two keyframes, rather than all video frames.
+ControlVideo [62] extends ControlNet [61] with fully cross-frame attention to joint edit all frames and further improves the performance with interleaved-frame smoother.
+TokenFlow [19] enhance PNP [51] with extended-attention to jointly edit a few keyframes at each denoising step and propagate the edit frames to the entire video.
+However, conducting hundreds of denoising steps with extended attention across all frames is time-consuming.
+In contrast, we only utilize diffusion models to edit a few keyframes, rather than all frames.
 Video Frame Interpolation (VFI).
 VFI aims to generate intermediate images between a pair of frames, which can be applied to creating slow-motion videos and enhancing refresh rate.
-Advanced methods typically entail estimating dense motions between frames, like optical flow, and subsequently warping the provided frames to generate intermediate ones [31, 23, 39, 28, 36, 22].
-However, these methods are most effective with simple or monotonous motion.
-Thus, they cannot be directly applied to the second stage.
-In our work, we perform frame interpolation by incorporating additional structural signals.
+Advanced methods typically estimate dense motions between frames, like optical flow, and subsequently warping the provided frames to generate intermediate ones [37, 27, 46, 34, 43, 26].
+However, these regression methods are most effective with simple or monotonous motion.
+In our work, we design a generative model to perform frame interpolation with additional structural signals, enabling interpolation with large motion.
 3 Preliminaries
@@
 In the first stage, an image is quantized into a sequence of discrete tokens via a Vector-Quantized (VQ) auto-encoder [16].
-In detail, given an image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, an encoder $\mathcal{E}$ encodes it into a series of latent vectors and discretize them through a nearest neighbour look up in a codebook of quantized embeddings with size $M$.
-To this end, an image can be represented with a sequence of codebook’s indices $\mathbf{Z}=[z_{i}]_{i=1}^{h\times w},z_{i}\in\{1,2,...,M\}$, where $h$ and $w$ is the resolution of latent features.
+In detail, given an image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, an encoder $\mathcal{E}$ encodes it into a series of latent vectors, which are then discretized via nearest neighbor lookup in a codebook of quantized embeddings with size $M$.
+Thus, an image can be represented with a sequence of codebook’s indices $\mathbf{Z}=[z_{i}]_{i=1}^{h\times w},z_{i}\in\{1,2,...,M\}$, where $h$ and $w$ is the resolution of latent features.
 A decoder $\mathcal{D}$ can reconstruct the indices back to image $\mathcal{D}(\mathbf{Z})\approx\mathbf{I}$.
@@
 $\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\mathbf{Z}\in\mathbb{D}}\left[\sum_{\bar{z_{i}}=\texttt{[MASK]}}-\log p_{\Phi}(z_{i}|\mathbf{\bar{Z}},\mathbf{c})\right].$ (1)
-During inference time, non-autoregressive decoding is applied to generate images.
-Specifically, given the conditions $\mathbf{c}$, all tokens are initialized as [MASK] tokens.
+During inference, non-autoregressive decoding is applied to generate images.
+Specifically, all tokens of an image or video are initialized as [MASK] tokens.
 At step $k$, all tokens are predicted in parallel while only tokens with the highest prediction scores are kept.
@@
 Latent Diffusion Models
-Denoising Diffusion Probabilistic Models (DDPM) [18] generate images through a progressive noise removal process applied to an initial Gaussian noisy image, carried out over a span of $T(>>K)$ time steps.
-To enable efficient high-resolution image generation, Latent Diffusion models [37] operates the diffusion process in the latent space of an autoencoder.
-First, an encoder $\mathcal{E^{\prime}}$ compresses an image $\mathbf{I}$ to a low-resolution latent code $x=\mathcal{E^{\prime}}(\mathbf{I})\in\mathbb{R}^{h\times w\times c}$.
-Second, a U-Net $\epsilon_{\theta}$ with attention modules [44] is trained to remove the noise with loss function:
-$\mathcal{L}_{LDM}=\mathbb{E}_{x_{0},\epsilon\sim N(0,I),t\sim}\|\epsilon-\epsilon_{\theta}(x_{t},t,\tau)\|_{2}^{2},$ (2)
-where $\tau$ is the text prompt and $x_{t}$ is the noisy latent sample at timestep $t$.
-Stable Diffusion is trained on datasets with billion scale text-image pairs, which serves as the foundation model of many generation tasks.
+Denoising Diffusion Probabilistic Models [22] generate images through a progressive noise removal process applied to an initial Gaussian noisy image, carried out over a span of $T$ time steps.
+To enable efficient high-resolution image generation, LDM [44] operates the diffusion process in the latent space of an autoencoder.
+In detail, an encoder $\mathcal{E^{\prime}}$ first compresses an image $\mathbf{I}$ to a low-resolution latent code $x=\mathcal{E^{\prime}}(\mathbf{I})\in\mathbb{R}^{h\times w\times c}$.
+Second, a U-Net $\epsilon_{\theta}$ with attention modules [53] is trained to remove the noise with loss
+$\mathcal{L}_{LDM}=\mathbb{E}_{x_{0},\epsilon\sim N(0,I),t\sim}\|\epsilon-\epsilon_{\theta}(x_{t},t,\tau)\|_{2}^{2}$ (2)
+, where $\tau$ is the text prompt and $x_{t}$ is the noisy latent sample at timestep $t$.
 ControlNet
-In practice, it’s challenging to use text prompt to describe the layout of generated image.
-Furthermore, ControlNet [52] is proposed to provide spatial layout conditions such as edge map, depth map, and human poses.
-In detail, ControlNet train the same U-Net architecture as SD and finetune it with specific conditions.
-We denote ControlNet as $\epsilon_{\theta}^{\prime}(x_{t},t,\tau,s)$, where $s$ is the spatial layout condition.
+Since accurately describing the layout of generated images using only text prompts is challenging, ControlNet [61] extends LDM by incorporating spatial layout conditions such as edge maps, depth maps, and human poses.
+In detail, ControlNet train the same U-Net architecture as LDM and finetune it with additional conditions.
+We denote it as $\epsilon_{\theta}^{\prime}(x_{t},t,\tau,s)$, where $s$ is the spatial layout.
 Figure 2: Overview of MaskINT. MaskINT disentangle the video editing task into two separate stages, i.e., keyframes joint editing and structure-aware frame interpolation.
@@
 Fig.
-2 shows an overview of our framework.
-We disentangle the video editing task into keyframe joint editing stage and structure-aware frame interpolation stage.
-Specifically, given a video clip with $N$ frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$, in the first stage, with the input text prompt $\mathbf{\tau}$, we simultaneously edit two keyframes, i.e., the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$, using existing image-editing model $g(\cdot)$ that requires no additional tuning.
-This frame based joint editing module provides high-quality coherent edited frames $\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N}=g(\mathbf{I}^{1},\mathbf{I}^{N},\mathbf{\tau})$ and is highly efficient on a pair of frame.
-In the second stage, we propose MaskINT to perform structure-aware frame interpolation via non-autoregressive transformers.
+2 shows an overview of our MaskINT.
+MaskINT disentangle the video editing task into keyframe editing stage and structure-aware frame interpolation stage.
+Specifically, given a video clip with $N$ frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$, in the first stage, with the input text prompt $\mathbf{\tau}$, we simultaneously edit two keyframes, i.e., the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$, using existing image-editing model $g(\cdot)$ in a zero-shot way.
+This process results in the creation of high-quality coherent edited frames $\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N}=g(\mathbf{I}^{1},\mathbf{I}^{N},\mathbf{\tau})$.
+In the second stage, MaskINT performs structure-aware frame interpolation via non-autoregressive transformers.
 The entire edited video frames are generated by $\{\mathbf{{}^{*}I}^{i}\}_{i=1}^{N}=f_{\Phi}(\mathbf{{}^{*}I}^{1},\mathbf{{}^{*}I}^{N},\{\mathbf{S}^{i}\}_{i=1}^{N})$, where $\mathbf{S}^{i}\in[0,1]^{H\times W\times 1}$ is the structural condition (i.e., the HED edge map [57]).
-Our MaskINT is trained with masked token modeling (MTM) on video only datasets, conditioning on the structural signal $\{\mathbf{S}^{i}\}_{i=1}^{N}$ as well as the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$.
+The structure-aware interpolation module is trained with masked token modeling (MTM) on video only datasets, conditioning on the structural signal $\{\mathbf{S}^{i}\}_{i=1}^{N}$ as well as the initial frame $\mathbf{I}^{1}$ and last frame $\mathbf{I}^{N}$.
 4.2 Keyframes Joint Editing
-To maintain the structure layout of selected keyframes, we take ControlNet [52] to edit $\mathbf{I}^{1}$ and $\mathbf{I}^{N}$ based on text prompt $\tau$ as well as their edge maps $\mathbf{S}^{1}$ and $\mathbf{S}^{N}$.
-However, even with the identical noise, applying ControlNet to each keyframes individually, i.e., $\epsilon_{\theta}^{\prime}(x_{t}^{1},t,\tau,\mathbf{S}^{1})$ and $\epsilon_{\theta}^{\prime}(x_{t}^{N},t,\tau,\mathbf{S}^{N})$, cannot guarantee the appearance consistency.
-To address this issue, following previous work [47, 15, 53], we extend the self-attention blocks to simultaneously process two keyframes.
-In detail, the self-attention block projects the noisy feature map $x_{t}^{j}$ of $j^{th}$ frame at time step $t$ into query $\mathbf{Q}^{j}$, key $\mathbf{K}^{j}$, value $\mathbf{V}^{j}$ in the original U-Net of SD.
-We extend the self-attention block to perform attention across all selected keyframes by concatenating their keys and values and calculate the attention by
-$\text{Softmax}(\dfrac{\mathbf{Q}^{j}[\mathbf{K}^{1},\mathbf{K}^{N}]^{\mathbf{T}}}{\sqrt{c}})[\mathbf{V}^{1},\mathbf{V}^{N}]$ (3)
-Note that although only two frames were used in Eq.3, this joint editing can seamlessly generalize to any number of frames, but at the cost of longer processing time and huge demand in resources.
-4.3 MaskINT
+To maintain the structure layout, we take ControlNet to edit $\mathbf{I}^{1}$ and $\mathbf{I}^{N}$ based on text prompt $\tau$ as well as their edge maps $\mathbf{S}^{1}$ and $\mathbf{S}^{N}$.
+However, even with the identical noise, applying ControlNet to each keyframes individually cannot guarantee the appearance consistency.
+To address this issue, following previous work [56, 19, 62], we extend the self-attention to simultaneously process two keyframes.
+In detail, the self-attention module projects the noisy feature map $x_{t}^{j}$ of $j^{th}$ frame at time step $t$ into query $\mathbf{Q}^{j}$, key $\mathbf{K}^{j}$, value $\mathbf{V}^{j}$ in the original U-Net.
+We extend the self-attention block to perform attention across all selected keyframes by concatenating their keys and values:
+$\text{Softmax}(\dfrac{\mathbf{Q}^{j}[\mathbf{K}^{1},\mathbf{K}^{N}]^{\mathbf{T}}}{\sqrt{c}})[\mathbf{V}^{1},\mathbf{V}^{N}].$ (3)
+Since each frame queries all frames and aggregates information from all of them, the appearance consistency can be achieved.
+Although only two frames were used in Eq.
+3, this joint editing can be seamlessly extended to any number of frames, but at the cost of increased processing time and significant GPU memory demand.
+4.3 Structure-Aware Frame Interpolation
 Structural-aware embeddings.
-Previous works [16, 51] demonstrate that non-autoregressive masked generative transformers can effectively perform frame prediction and interpolation tasks.
-However, these architectures lack explicit control of structure, making it difficult to follow the motion of original videos.
-In our work, we explicitly introduce structural condition of each frame into the generation process.
-Specifically, we tokenize both RGB frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$ and structure maps $\{\mathbf{S}^{i}\}_{i=1}^{N}$ with an off-the-shelf 2D VQ tokenizer [12].
-We utilize 2D VQ rather than 3D VQ [51] to accommodate varying numbers of frames and frame rate without constraints.
-We denote the tokens from RGB frames as $\mathbf{{}^{c}Z}=\{^{c}z_{i}\}_{i=1}^{N\times h\times w}$ (color token) and tokens from edge maps as $\mathbf{{}^{s}Z}=\{^{s}z_{i}\}_{i=1}^{N\times h\times w}$ (structure token), where ${}^{s}z_{i}$, ${}^{c}z_{i}\in\{1,2,...,M\}$, where $M$ is the codebook size.
-Subsequently, two distinct embedding layers $e^{C}(\cdot)$ and $e^{S}(\cdot)$ are learned to map token indices $\mathbf{{}^{c}Z}$ and $\mathbf{{}^{s}Z}$ into their respective embedding spaces.
+Non-autoregressive masked generative transformers [20, 60] can effectively perform frame prediction and interpolation tasks.
+However, these works lack explicit control of structure, making it difficult to follow the motion of input videos.
+Thus, we explicitly introduce structural condition of each frame into the generation process.
+In detail, we tokenize both RGB frames $\{\mathbf{I}^{i}\}_{i=1}^{N}$ and structure maps $\{\mathbf{S}^{i}\}_{i=1}^{N}$ with an off-the-shelf 2D VQ tokenizer [16].
+We utilize 2D VQ rather than 3D VQ [60] to accommodate diverse numbers of frames and frame rate.
+We denote tokens from RGB frames as $\mathbf{{}^{c}Z}=\{^{c}z_{i}\}_{i=1}^{N\times h\times w}$ (color token) and tokens from edge maps as $\mathbf{{}^{s}Z}=\{^{s}z_{i}\}_{i=1}^{N\times h\times w}$ (structure token), where ${}^{s}z_{i}$, ${}^{c}z_{i}\in\{1,2,...,M\}$ and $M$ is the codebook size.
+Subsequently, two distinct embedding layers $e^{C}(\cdot)$ and $e^{S}(\cdot)$ are learned to map tokens $\mathbf{{}^{c}Z}$ and $\mathbf{{}^{s}Z}$ into their respective embedding spaces.
 Learnable 2D spatial positional encoding $\mathbf{P^{S}}$ and temporal positional encoding $\mathbf{P^{T}}$ are also added [4].
@@
 Transformer with Window-Restricted Attention.
-Previous masked generative transformers [7, 51] employ a pure transformer with global attention [44].
-However, Given that there is no substantial motion between consecutive frames, we adopt self-attention within a restricted window to further mitigate computational overhead, following MaskViT [16].
-In detail, our approach involves two distinct stages of attention.
-Initially, we employ spatial window attention, confining attention to tokens within a frame of dimensions $1\times h\times w$.
-Subsequently, we extend this to spatial-temporal window attention, which confines attention to tokens within a tube of dimensions $N\times h_{w}\times w_{w}$, where $h_{w}$ and $w_{w}$ is the window size, where $h_{w}<h$ and $w_{w}<w$, which greatly reduce the complexity.
+Previous masked generative transformers [10, 60] employ the standard transformer with global attention [53].
+Given that there is no substantial motion between consecutive frames, we adopt self-attention within a restricted window [20] to further mitigate computational overhead In detail, our approach involves two distinct stages of attention.
+First, we conduct spatial window attention, where attention is calculated within each frame of dimensions $1\times h\times w$.
+We then conduct spatial-temporal window attention, where attention is calculated within a tube of dimensions $N\times h_{w}\times w_{w}$, where $h_{w}$ and $w_{w}$ is the window size.
 Besides, to further reduce computation of transformer, we also add a shallow convolution layers to downsample $\mathbf{X}$ before the transformer encoder layers and an upsample layer at the end.
 Training.
-By fully disentangling keyframes editing and frame interpolation into distinct stages, our model no longer necessitates paired videos for training.
-Consequently, we can train MaskINT using video only datasets.
+We train MaskINT using video-only datasets.
 Denote the color token of $i^{th}$ RGB frame as $\mathbf{{}^{c}Z}^{i}=\{^{c}z^{i}\}_{i=1}^{h\times w}$.
@@
 The structure-aware window-restricted transformer with parameters $\Theta$ is trained by
-$\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\mathbf{{}^{c}Z},\mathbf{{}^{s}Z}\in\mathbb{D}}[\sum_{\bar{{}^{c}z_{i}}=\texttt{[MASK]}}-\log p_{\Theta}(^{c}z_{i}|\mathbf{\bar{{}^{c}Z}},\mathbf{{}^{s}Z})]$ (4)
+$\mathcal{L}_{MTM}=\mathop{\mathbb{E}}_{\{\mathbf{{}^{c}Z},\mathbf{{}^{s}Z}\}\in\mathbb{D}}[\sum_{\bar{{}^{c}z_{i}}=\texttt{[MASK]}}-\log p_{\Theta}(^{c}z_{i}|\mathbf{\bar{{}^{c}Z}},\mathbf{{}^{s}Z})].$ (4)
 Inference.
-During the inference time, our MaskINT can seamlessly generalize to perform frame interpolation between the jointly edited frames, although it is only trained with regular videos.
+During inference, MaskINT can seamlessly generalize to perform frame interpolation between the jointly edited frames, although it is only trained with regular videos.
 Specifically, we tokenize the the initial and last edited frames $\mathbf{{}^{*}I}^{1}$ and $\mathbf{{}^{*}I}^{N}$ from Stage 1 into color tokens $\mathbf{{}^{c}_{*}Z}^{1}$ and $\mathbf{{}^{c}_{*}Z}^{N}$, and initialize color tokens of all intermediate frames $\{\mathbf{\bar{{}^{c}Z}}^{2},...,\mathbf{\bar{{}^{c}Z}}^{N-1}\}$ with [MASK] tokens.
@@
 Implementation Details.
-We train our model with $100k$ videos from ShutterStock website [^1] https://www.shutterstock.com/video.
-During training time, we random select a $T=16$ video clip with frame interval $1,2,4$ from each video and resize it to $384\times 672$.
+We train our model with $100k$ videos from ShutterStock [1].
+During training, we random select a video clip with $T=16$ frames and frame interval of $1,2,4$ from each video.
+All frames are resized to $384\times 672$.
 We utilize Retina-VQ [15] with $8$ downsample ratio, i.e., each frame has $48\times 84$ tokens.
@@
 The initial learning rate is set to $41e-$ 4 and decayed with cosine schedule.
-During the inference time, we set the number of decoding step $K$ to $32$ and the temperature $t$ to $4.5$.
+We set the number of decoding step $K$ to $32$ and the temperature $t$ to $4.5$ in inference.
 Evaluation.
-Following [47], we use the selected $40$ object-centric videos of the DAVIS dataset [33], covering humans, animals, vehicles, etc.
-Besides, we also select $30$ unseen videos from the ShutterStock dataset.
+We evaluate our method on $40$ selected object-centric videos [56] from the DAVIS dataset [40], as well as $30$ unseen videos from the ShutterStock [1], covering animals, vehicles, etc.
 For each video, we manually design $5$ edited prompts, including object editing, background changes and style transfers.
-Following previous works [53, 34], we assess the quality of the generated videos using CLIP [35].
-In detail, we evaluate 1) Temporal consistency, which calculates the average cosine similarity of all pairs of consecutive frames.
-2) Prompt consistency, which calculates the average cosine similarity between given text prompt and all video frames.
-To evaluate the efficiency, we report the duration required for generating a 16-frame video clip on a single NVIDIA A6000 GPU.
+Following previous works [62, 41, 19, 8], we evaluate the quality of the generated videos with the following metrics: 1) Prompt consistency, which calculates the CLIP [42] embedding similarity between text prompt and all video frames.
+2) Temporal consistency, which calculates the CLIP embedding similarity of all pairs of consecutive frames.
+3) Warpping-error [30], which computes the optical flow between consecutive frames using RAFT [50], and warps the edited frames to the next according it.
+4) Long-term temporal consistency.
+Inspired by the global coherence metric in [52], we also calculate the warp error between all possible pairs of frames as the long-term consistency.
+As for efficiency, we report the running time for generating a 16-frame video clip on a single NVIDIA A6000 GPU.
 Figure 3: Examples of video editing with MaskINT. Frames with red bounding box are jointly edited keyeframes.
 5.2 Results
-We select methods that built upon text-to-image diffusion models for comparison, including TokenFlow [15], Text-to-video zero [25], and ControlVideo [53].
-We also consider apply ControlNet to each frame individually with the same initial noise as baseline.
-Besides, we also compare with frame interpolation with FILM [36] with the same edited keyframes.
+We apply ControlNet to each frame individually with the same noise as the baseline.
+We select methods that built upon T2I diffusion models for comparison, including Tune-A-Video [56], TokenFlow [19], Text-to-video zero [29], and ControlVideo [62].
+Besides, we compare with VFI methods such as FILM [43] using the same edited keyframes.
 Method DAVIS ShutterStock Time
-T.C. P.C. T.C. P.C
-ControlNet per frame [52] 0.9137 0.3136 0.9419 0.3040 50s
-Text2Video-zero [25] 0.9642 0.3124 0.9811 0.3036 60s
-TokenFlow [15] 0.9774 0.3169 0.9869 0.3133 150s
-ControlVideo-edge [53] 0.9746 0.3143 0.9864 0.3032 120s
-MaskINT (ours) 0.9519 0.3112 0.9714 0.3038 22s
-Table 1: Quantitative comparisons on the quality of the generated videos. “T.C.” stands for “temporal consistency”, and “P.C.” stands for “prompt consistency”.
+[…]
+/×$\times 10^{-3}$ /×$\times 10^{-3}$ /×$\times 10^{-3}$ /×$\times 10^{-3}$
+ControlNet per frame [61] 0.314 0.914 36.6 58.6 0.304 0.942 25.9 45.1 50s
+Tune-a-Video [56] 0.299 0.966 20.4 48.2 0.292 0.979 13.4 34.9 20min
+Text2Video-zero [29] 0.312 0.964 20.7 42.5 0.304 0.981 16.0 33.0 60s
+ControlVideo-edge [62] 0.314 0.975 6.9 23.9 0.303 0.986 7.4 20.5 120s
+TokenFlow [19] 0.317 0.977 7.0 19.6 0.313 0.987 5.4 15.8 150s
+MaskINT (ours) 0.311 0.952 9.5 27.7 0.304 0.971 8.6 22.3 22s
+Table 1: Quantitative comparisons.
+“T.C." stands for “temporal consistency", “P.C." stands for “prompt consistency", “W.E" stands for “warpping-error", and “L.C." stands for “long-term temporal consistency".
 Quantitative Comparisons.
-Table 1 summarize the performance of these methods on both DAVIS and ShutterStock datasets.
-Notably, our method achieves comparable performance with diffusion methods, in terms of both temporal consistency and prompt consistency, while brings a significant acceleration in processing speed.
-In detail, MaskINT is almost 5.5 times faster than ControlVideo [53], whose fully cross-frame attention is computationally extensive.
+Tab.
+1 summarize the performance of these methods on both DAVIS and ShutterStock datasets.
+Notably, our method achieves comparable performance with diffusion methods, in terms of all four evaluation metrics, while brings a significant acceleration in processing speed.
+In detail, the test-time finetuning of TAV [56] is extremely time-consuming, making it a less practical solution.
+MaskINT is almost 5.5 times faster than ControlVideo [62], whose fully cross-frame attention is computationally extensive.
 Moreover, MaskINT is nearly 7 times faster than TokenFlow [19], whose DDIM inversion is time-consuming.
-On the contrary, our acceleration is derived from a combination of a lightweight network design and a reduced number of decoding steps in masked generative transformers.
+The efficiency of our method is derived from a combination of a lightweight network and a reduced number of decoding steps with masked generative transformers.
 Figure 4: Qualitative comparisons with diffusion-based methods. More examples are shown in Supplementary.
@@
 1 and Fig.
-3 show several samples of edited videos with MaskINT.
-Our method is capable of generating temporally consistent videos that adhere to the provided text prompts.
-This extends to a wide range of applications with text prompts, encompassing tasks such as stylization, background editing, foreground editing, and more.
-It also works well on challenging videos with substantial motion, such as jumping and running.
-Moreover, Fig.
-4 provides a qualitative comparison of MaskINT to other baselines.
-Remarkably, diffusion methods [15, 25, 53] can ensure the consistency of overall appearance, but sometimes cannot maintain the consistency of detailed regions.
-For example, both TokenFlow [15] and Text2Video-Zero[25] exhibit noticeable artifacts in the leg region of the human subjects.
-ControlVideo [53] produces inconsistent hats.
+3 show edited video samples using MaskINT.
+Our method is able to generate temporally consistent videos in accordance with text prompts, enabling diverse applications including stylization, background and foreground editing, and beyond.
+It is also effective in videos with complex motion, such as jumping and running, as well as scenarios with multiple subjects.
+Besides, although the structure-aware interpolation module is trained with realistic videos only, it still generalizes to diverse synthesized images.
+For example, we can successfully generate a video in the style of Van Gogh even in the absence of videos for this particular style in the training set.
+Fig.
+4 provides a comparison of MaskINT to other methods.
+Remarkably, diffusion-based methods with extended attention [56, 19, 29, 62] can maintain the overall appearance consistency, but sometimes result in inconsistent details.
+For example, TokenFlow [19] and Text2Video-Zero[29] exhibit noticeable artifacts in the leg region of the human subjects and ControlVideo [62] produces inconsistent hats.
 The potential explanation lies in the fact that these methods offer control over temporal consistency implicitly.
-Furthermore, FILM [36] produces videos that deviate from the original motions with the same edited keyframes.
+Besides, FILM [43] generates videos that cannot follow the original motions.
 Our MaskINT consistently interpolates the intermediate frames based on the structure condition and even maintain better consistency in local regions.
 Extension on Long Video Editing.
-Since the non-autoregressive pipeline generates all video frames simultaneously, it’s challenging for it to edit an entire long video due to GPU memory limitation.
-Nevertheless, our framework can still be extended to generate long videos by dividing the long video into short clips and progressively performing frame interpolation within each clip.
+Given that the non-autoregressive pipeline generates all frames simultaneously, it’s challenging to edit an entire long video at once due to GPU memory limitation.
+Nevertheless, our framework can be extended to handle long videos by dividing them into short clips and progressively performing frame interpolation within each segment.
 For instance, given a video with $60$ frames, we select the $1^{\text{st}}$, $16^{\text{th}}$, $31^{\text{st}}$, $46^{\text{th}}$, and $60^{\text{th}}$ frames as keyframes.
@@
 5, with this design, our method can still generate consistent long videos.
-Besides, in this proposed extension, the generation of later frames is decoupled from the generated early frames, which differs from the autoregressive long-video generation pipeline in Phenaki [45].
+Moreover, in this proposed extension, the generation of later frames is dissociated from that of early frames, which differs from the autoregressive generation pipeline in Phenaki [54].
 Consequently, even if some early frames encounter difficulties, the generation of later frames can still proceed successfully.
@@
 5.3 Ablation Studies
-Video Frame Interpolations
-We conduct a quantitative evaluation on the performance of our structure-aware interpolation module.
-In this ablation study, we perform frame interpolation using the original keyframes to reconstruct the original intermediate frames, and compare the interpolated frames with the original video frames at the pixel level.
+Video Frame Interpolations.
+In this ablation study, we evaluate the performance of the structure-aware interpolation module.
+In detail, we perform frame interpolation using original keyframes, and compare the interpolated frames with the original intermediate video frames.
 We use same testing samples from DAVIS and Shutterstock datasets, employing peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and structured similarity (SSIM) as the evaluation metrics.
@@
 We also showcase the performance of applying VQ-GAN [16] to all video frames, serving as an upper bound for our method.
-As in Table 2, our method significantly outperforms VFI methods on all evaluation metrics, with the benefit of the structure guidance from the intermediate frames.
+As in Tab.
+2, our method significantly outperforms VFI methods on all evaluation metrics, with the benefit of the structure guidance from the intermediate frames.
 Furthermore, Fig.
-6 shows qualitative comparison between video frame interpolation methods FILM [36] and MaskINT.
-Even when confronted with significant motion between two frames, our approach successfully reconstructs the original video, maintaining consistent motion through the aid of structural guidance.
+6 shows qualitative comparison between FILM [43] and MaskINT.
+Even when confronted with significant motion between two frames, our method successfully reconstructs the original video, maintaining consistent motion through the aid of structural guidance.
 In contrast, FILM introduces undesirable artifacts, including disorted background, duplicated cat hands, and the absence of a camel’s head, etc.
-The major reason is that current VFI models mainly focus on generating slow-motion effects and enhancing frame rate, making them less effective in handling frames with large motions, which usually requires a better semantic understanding.
-Additionally, the absence of structural guidance poses a challenge for these VFI methods in accurately aligning generated videos with the original motion.
+The major reason is that current VFI models mainly focus on generating slow-motion effects and enhancing frame rate, making them less effective in handling frames with complex motions, which usually requires a better semantic understanding.
+Besides, the absence of structural guidance poses a challenge for them in accurately aligning generated videos with the original motion.
 Figure 6: Qualitative comparisons on video reconstruction with original RGB frames. Frames with red bounding box are given.
@@
 FILM [43] 17.00 0.5011 0.2363 20.90 0.7453 0.1246
-MaskINT 22.15 0.6332 0.1483 24.19 0.7616 0.1097
+Ours 22.15 0.6332 0.1483 24.19 0.7616 0.1097
 VQGAN (ground truth) 25.66 0.7429 0.0784 27.81 0.8327 0.0561
@@
 Generally, with more information, performing frame interpolation is easier.
-However, simultaneously editing more keyframes requires longer time due to the global attention among them.
+However, simultaneously editing more frames requires longer time due to the global attention among them.
 Decoding steps
-We also explore the number of decoding steps of the masked generative transformers in the second stage.
-The right part of Table 3 shows that more decoding steps can bring slight improvement on the temporal consistency, but requires more time.
+We also explore the number of decoding steps $K$ for the masked generative transformers in the second stage.
+The right part of Tab.
+3 shows that more decoding steps can bring slight improvement on the temporal consistency, but requires more time.
 Considering the trade-off between performance and efficiency, we chose $K=32$ steps by default in all experiments.
-# keyframes T.C. P.C. Time 1 0.9690 0.2984 19s 2 0.9714 0.3038 22s 3 0.9721 0.3051 26s 4 0.9728 0.3069 29s 6 0.9737 0.3035 35s
-#decoding step $K$ T.C. P.C. Time 16 0.9691 0.3038 15s 32 0.9714 0.3038 22s 64 0.9719 0.3040 33s 128 0.9720 0.3041 62s
-Table 3: Ablation study on ShutterStock dataset of the number of keyframes and the number of decoding steps $K$. “T.C.” stands for “temporal consistency” and “P.C.” stands for “prompt consistency”.
+# keyframes T.C. P.C. Time
+1 0.9690 0.2984 19s
+2 0.9714 0.3038 22s
+3 0.9721 0.3051 26s
+4 0.9728 0.3069 29s
+6 0.9737 0.3035 35s
+#decoding step $K$ T.C. P.C. Time
+16 0.9691 0.3038 15s
+32 0.9714 0.3038 22s
+64 0.9719 0.3040 33s
+128 0.9720 0.3041 62s
+Table 3: Ablation study on the number of keyframes and the number of decoding steps $K$. “T.C." stands for “temporal consistency" and “P.C." stands for “prompt consistency".
 6 Limitation and Future Work
-One potential limitation of our approach is that it still necessitates the use of diffusion-based methods for keyframe editing in the initial stage, which can still be somewhat time-consuming.
-As for the future work, given that our approach disentangles video editing into two distinct stages, we intend to explore the integration of token-based methods, like Muse [8], for image editing.
+One limitation of MaskINT is that it can only perform structure-preserving video editing, such as altering style or appearance.
+Thus, it cannot handle edits that require structural changes, for example, change a dog to a horse, a limitation shared with TokenFlow [19].
+It also requires that no new objects should appear in the intermediate frames.
+Besides, the performance of MaskINT highly relies on the tedious image-editing model and structure detector.
+When these models fail, the structure-aware interpolation becomes meaningless, resulting in artifacts.
+In the future, given that our approach disentangles video editing into two distinct stages, we intend to explore the integration of token-based methods, like Muse [11], for image editing.
 This endeavor aims to further enhance the efficiency of the initial stage.
 7 Conclusion
-We propose MaskINT towards consistent and efficient video editing with text prompt.
-MaskINT disentangle this task into keyframes joint editing with diffusion methods and structure-aware frame interpolation with non-autoregressive masked transformers.
-Experimental results demonstrate that MaskINT achieves comparable performance with pure diffusion-based methods while significantly reduces the inference time.
-Moreover, our work demonstrates the substantial promise of non-autoregressive generative transformers within the realm of video editing.
+We propose MaskINT towards consistent and efficient video editing with prompt.
+MaskINT disentangles this task into keyframes joint editing with T2I model and structure-aware frame interpolation with non-autoregressive masked transformers.
+Experimental results demonstrate that MaskINT achieves comparable performance with pure diffusion-based methods while significantly accelerate the inference up to 7x.
+We believe our work demonstrates the substantial promise of non-autoregressive generative transformers within the realm of video editing.
 Supplementary Material
+1 Dataset Details
+ShutterStock [1] is a commercial dataset with over 30 million paired text-video examples.
+Our model is trained using a subset of 100k videos.
+The WebVid dataset [2] offers an accessible version of ShutterStock, but they include a central watermark due to copyright constraints.
 2 Diverse Structural Conditions
@@
 Table 1: Quantitative comparisons of the combination of varied structural conditions in each stage on ShutterStock.
-“T.C.” stands for “temporal consistency”, and “P.C.” stands for “prompt consistency”.
+“T.C." stands for “temporal consistency", and “P.C." stands for “prompt consistency".
 3 Additional Examples for Comparisons